\begin{thebibliography}{10}

\bibitem{Arkhangelskii_Ponomarev}
A.~V. Arkhangelskii and V.~I. Ponomarev.
\newblock {\em {O}snovy obshchei topologii v zadachakh i uprazhneniyakh.}
\newblock Izdat. ``Nauka'', Moscow, 1974.

\bibitem{Arnautov_book}
V.~I. Arnautov, S.~T. Glavatsky, and A.~V. Mikhalev.
\newblock {\em Introduction to the theory of topological rings and modules},
  volume 197 of {\em Monographs and Textbooks in Pure and Applied Mathematics}.
\newblock Marcel Dekker, Inc., New York, 1996.

\bibitem{Arnautov_Ursul}
V.~I. Arnautov and M.~I. Ursul.
\newblock Uniqueness of a linearly compact topology in rings.
\newblock {\em Mat. Issled.}, (53):6--14, 221, 1979.
\newblock Topological spaces and algebraic systems.

\bibitem{Bourbaki_II}
N.~Bourbaki.
\newblock {\em { {K}ommutativnaya algebra.}}
\newblock Izdat. ``Mir'', Moscow, 1971.
\newblock {{{\`E}}lementy Matematiki, Vyp.} XXVII, XXVIII, XXX, XXXI.
  [Foundations of Mathematics, No. XXVII, XXVIII, XXX, XXXI], Translated from
  the French by A. A. Belski\u\i , Edited by E. S. Golod.

\bibitem{Bourbaki}
N.~Bourbaki.
\newblock {\em { {O}bshchaya topologiya.}}
\newblock Izdat. ``Nauka'', Moscow, 1975.
\newblock {{{I}}spol$\prime$zovanie veshchestvennykh chisel v obshche\u\i
  topologii. Funktsional$\prime$nye prostranstva. Svodka rezul$\prime$tatov.
  Slovar$\prime$.} [Application of real numbers in general topology. Functional
  spaces. Resum\'e of results. Vocabulary], Translated from the third French
  edition by S. N. Kra\v ckovski\u\i , Edited by D. A. Ra\u\i kov.

\bibitem{Bourbaki_III}
N.~Bourbaki.
\newblock {\em General topology. {C}hapters 1--4}.
\newblock Elements of Mathematics (Berlin). Springer-Verlag, Berlin, 1998.
\newblock Translated from the French, Reprint of the 1989 English translation.

\bibitem{Dikranjan}
D.~Dikranjan.
\newblock Minimal topological rings.
\newblock {\em Serdica}, 8(2):149--165, 1982.

\bibitem{Engelking}
R.~Engelking.
\newblock {\em General topology}, volume~6 of {\em Sigma Series in Pure
  Mathematics}.
\newblock Heldermann Verlag, Berlin, ii edition, 1989.
\newblock Translated from the Polish by the author.

\bibitem{Freudenthal}
H.~Freudenthal.
\newblock Einige {S}\"atze \"uber topologische {G}ruppen.
\newblock {\em Ann. of Math. (2)}, 37(1):46--56, 1936.

\bibitem{Gaughan}
E.~D. Gaughan.
\newblock Topological group structures of infinite symmetric groups.
\newblock {\em Proc. Nat. Acad. Sci. U.S.A.}, 58:907--910, 1967.

\bibitem{Graev}
M.~I. Graev.
\newblock Theory of topological groups. {I}. {N}orms and metrics on groups.
  {C}omplete groups. {F}ree topological groups.
\newblock {\em Uspehi Matem. Nauk (N.S.)}, 5(2(36)):3--56, 1950.

\bibitem{Hewitt_Ross}
E.~Hewitt and K.~A. Ross.
\newblock {\em Abstract harmonic analysis. {V}ol. {I}}, volume 115 of {\em
  Grundlehren der Mathematischen Wissenschaften [Fundamental Principles of
  Mathematical Sciences]}.
\newblock Springer-Verlag, Berlin-New York, second edition, 1979.
\newblock Structure of topological groups, integration theory, group
  representations.

\bibitem{Hochster_Kiltinen}
M.~Hochster and J.~O. Kiltinen.
\newblock Commutative rings with identity have ring topologies.
\newblock {\em Bull. Amer. Math. Soc.}, 76:419--420, 1970.

\bibitem{Hulanicki}
A.~Hulanicki.
\newblock On locally compact topological groups of power of continuum.
\newblock {\em Fund. Math.}, 44:156--158, 1957.

\bibitem{Jacobson_2}
N.~Jacobson.
\newblock Totally {D}isconnected {L}ocally {C}ompact {R}ings.
\newblock {\em Amer. J. Math.}, 58(2):433--449, 1936.

\bibitem{Jacobson_3}
N.~Jacobson.
\newblock A {N}ote on {T}opological {F}ields.
\newblock {\em Amer. J. Math.}, 59(4):889--894, 1937.

\bibitem{Jacobson}
N.~Jacobson.
\newblock {\em Structure of rings}.
\newblock American Mathematical Society, Colloquium Publications, vol. 37.
  American Mathematical Society, 190 Hope Street, Prov., R. I., 1956.

\bibitem{Jones}
F.~B. Jones.
\newblock On the first countability axiom for locally compact {H}ausdorff
  spaces.
\newblock {\em Colloq. Math.}, 7:33--34, 1959.

\bibitem{Kaplansky}
I.~Kaplansky.
\newblock Topological rings.
\newblock {\em Amer. J. Math.}, 69:153--183, 1947.

\bibitem{Kaplansky_selected}
I.~Kaplansky.
\newblock {\em Selected papers and other writings}.
\newblock Springer Collected Works in Mathematics. Springer, New York, 2013.
\newblock With an introduction by Hyman Bass, Reprint of the 1995 edition
  [MR1340874].

\bibitem{Lam}
T.~Y. Lam.
\newblock {\em A first course in noncommutative rings}, volume 131 of {\em
  Graduate Texts in Mathematics}.
\newblock Springer-Verlag, New York, 1991.

\bibitem{Leptin}
H.~Leptin.
\newblock Linear kompakte {M}oduln und {R}inge.
\newblock {\em Math. Z.}, 62:241--267, 1955.

\bibitem{Scottish_Book}
R.~D. Mauldin, editor.
\newblock {\em The {S}cottish {B}ook}.
\newblock Birkh\"auser/Springer, Cham, second edition, 2015.
\newblock Mathematics from the Scottish Caf\'e with selected problems from the
  new Scottish Book, Including selected papers presented at the Scottish Book
  Conference held at North Texas University, Denton, TX, May 1979.

\bibitem{Mutylin}
A.~F. Mutylin.
\newblock Completely simple commutative topological rings.
\newblock {\em Mat. Zametki}, 5:161--171, 1969.

\bibitem{Pontrjagin_book}
L.~S. Pontryagin.
\newblock {\em Continuous groups}.
\newblock ``Nauka'', Moscow, fourth edition, 1984.

\bibitem{Skornjakov}
L.~Skornjakov.
\newblock Einfache lokal bikompakte {R}inge.
\newblock {\em Math. Z.}, 87:241--251, 1965.

\bibitem{Ursul}
M.~Ursul.
\newblock {\em Topological rings satisfying compactness conditions}, volume 549
  of {\em Mathematics and its Applications}.
\newblock Kluwer Academic Publishers, Dordrecht, 2002.

\bibitem{Ware_Zelmanowitz}
R.~Ware and J.~Zelmanowitz.
\newblock Simple endomorphism rings.
\newblock {\em Amer. Math. Monthly}, 77:987--989, 1970.

\bibitem{Warner_book_II}
S.~Warner.
\newblock {\em Topological fields}, volume 157 of {\em North-Holland
  Mathematics Studies}.
\newblock North-Holland Publishing Co., Amsterdam, 1989.
\newblock Notas de Matem\'atica [Mathematical Notes], 126.

\bibitem{Warner_book}
S.~Warner.
\newblock {\em Topological rings}, volume 178 of {\em North-Holland Mathematics
  Studies}.
\newblock North-Holland Publishing Co., Amsterdam, 1993.

\bibitem{Zelinsky_I}
D.~Zelinsky.
\newblock Linearly compact modules and rings.
\newblock {\em Amer. J. Math.}, 75:79--90, 1953.

\end{thebibliography}

\documentclass[12pt,a4paper]{amsart}

\usepackage{amsfonts,amsmath,amsthm,  amssymb, amscd}

\newtheorem{theorem}{Theorem}
\newtheorem{corollary}{Corollary}

\newtheorem{lemma}{Lemma}

\newtheorem{remark}{Remark}

\keywords{Topological ring, Endomorphism ring, Bohr topology, Finite topology, Locally compact ring}
\subjclass{16W80; 16N20;  16S50; 16N40}
\thanks{
Supported by  UAEU grants:  UPAR  G00001922 and  GG00002160}

\title[Completely simple rings]
{Completely simple endomorphism rings \\of modules}
\author{V.~Bovdi, M.~Salim, Mihail Ursul}

\address{Department of Math. Sciences\\
UAE University\\
Al-Ain\\
United Arab Emirates}
\email{vbovdi@gmail.com}
\email{m.salim@uaeu.ac.ae}

\address{Department of Mathematics and Computer Science,\\ University of Technology, Lae,\\
Papua New Guinea}
\email{mihail.ursul@gmail.com}

\begin{document}
\maketitle
%\large

\begin{abstract}
It is proved that if $A_p$ is a countable elementary abelian $p$-group, then:
(i) The ring ${\rm End}\,(A_p)$ does not admit a nondiscrete locally compact ring topology.
(ii) Under (CH) the simple ring  ${\rm End}\,(A_p)/I $, where $I$ is the ideal of ${\rm End}\,(A_p)$ consisting of all endomorphisms with finite images,  does not admit a nondiscrete locally compact ring topology.
(iii) The finite topology  on  ${\rm End}\,(A_p)$ is the only second metrizable ring topology on it.
Moreover, a characterization of completely simple endomorphism rings of the endomorphism rings of modules over  commutative rings  is also obtained.
\end{abstract}

\section{Introduction}
The notion of associative simple ring can be extended for associative  topological rings in several  ways:
\begin{itemize}
\item[{(i)}]
simple abstract ring endowed with a nondiscrete ring topology (for example,  the classification of nondiscrete locally compact division rings, see \cite{Pontrjagin_book}, Chapter IV and  \cite{Bourbaki_II, Jacobson_2,  Jacobson_3}; we refer to  some historical notes about  locally compact division rings to  \cite{Warner_book_II});
\item[{(ii)}]
 topological ring without nontrivial closed ideals (for example,  see \cite{Leptin, Zelinsky_I}).
\item[{(iii)}]topological ring $R$ with the property that if $f:R\to S$ is a continuous homomorphism in a topological ring $S$, then either $f=0$ or $f$ is a topological embedding of $R$ into $S$ (see \cite{Mutylin}).
\end{itemize}

In all cases it is assumed that multiplication is not trivial.

I.~Kaplansky has mentioned  (see \cite{Kaplansky_selected}, p.56) that classification of locally compact simple rings in characteristic $p>0$  is difficult. He proved that every simple nondiscrete locally  compact simple torsion-free ring is a matrix ring over a locally  compact division  ring. However in \cite{Skornjakov} (see also \cite{Warner_book}) has been constructed a nondiscrete locally  compact simple ring of positive characteristic which is not a  matrix ring over a division ring. Thereby the program of classification of nondiscrete locally compact simple rings was finished. Nevertheless it is interesting  to look for new examples of locally compact simple rings.

If $A_p$ is a countable elementary abelian $p$-group and $I$ is an ideal of  ${\rm End}\,(A_p)$ consisting of endomorphisms with finite images, then the factor ring ${\rm End}\,(A_p)/I$ is a simple von Neumann regular ring. We prove that under (CH) this ring does not admit a nondiscrete locally compact ring topology.

\bigskip

S.~Ulam (see \cite{Scottish_Book}, Problem 96, p.181)  posed  the following problem: "{\it Can the group $S_\infty$ of all permutations of integers so metrized that the group operation (composition of permutations) is a continuous function and the set $S_\infty$ becomes, under this metric, a compact space? (locally compact?)}". E.D.~Gaughan (see \cite{Gaughan}) has solved this problem in the negative.

\bigskip

In  \S\ref{P:1}   we  study an analogous problem for  the endomorphism ring of a countable elementary abelian $p$-group, namely: "{\it Does the endomorphism ring ${\rm End}(A_p)$ of a countable elementary $p$-group $A_p$ admit a nondiscrete locally compact ring topology?}".  Similarly to Ulam's problem we obtain a negative answer to this question. Moreover, we prove that $\mathcal{T}_{fin}$ is the only ring topology $\mathcal{T}$ on ${\rm End}\,(A_p)$ such that $({\rm End}\,(A_p), \mathcal{T})$ is complete and second metrizable.

\bigskip

We classify  in \S4   the completely  simple rings $({\rm End}\,(M), \mathcal{T}_{fin})$  of  vector spaces $M$ over division rings.
Corollary \ref{C:2} gives a characterization of semisimple left linearly compact minimal rings.  It should be mentioned that Corollary \ref{C:2} is related to a result from \cite{Arnautov_Ursul} stating that any semisimple ring admits at most one linearly compact topology.

\bigskip

Furthermore,  we obtain in \S5 a description of  completely simple rings of the form $({\rm End}\,(M_R), \mathcal{T}_{fin})$ of modules $M$ over a commutative ring $R$. We extend the result of \cite{Ware_Zelmanowitz} to topological rings $({\rm End}\,(M_R), \mathcal{T}_{fin})$.

\section{Notation, Conventions and Preliminary Results}

Rings are assumed to be associative, not necessarily with identity. Topological spaces are assumed to be completely regular.
The {\it weight} (see \cite{Engelking}, p.12) of the space $X$ is denoted by $w(X)$. The {\it pseudocharacter} of a point $x\in X$  (see \cite{Engelking}, p.135) is the smallest cardinal of the form $|\mathcal{U}|$, where  $\mathcal{U}$  is a family of open subsets of $X$ such that $\cap\mathcal{U}=\{x\}$. The closure of a subset $A$ of the topological space $X$ is denoted by $\overline{A}$ and the interior by ${\rm Int}(A)$ (see \cite{Engelking}, p.14). A topological space $X$ is called a {\it Baire  space} (see \cite{Engelking}, p.198) if for each sequence $\{X_1,X_2,\ldots\}$ of open dense subsets of $X$ the intersection $\cap_{i=1}^\infty  G_i$ is a dense set.

An abelian group $A$ is called {\it elementary abelian $p$-group} ($p$ prime) if $pa=0$ for all $a\in A$.  It is well-known  that   $A$  is a direct sum of copies of the cyclic group $Z(p)$. The subring of a ring $R$ generated by a subset $S$, is denoted by $\langle S\rangle$. A ring $R$ is called {\it locally finite} if every its finite subset is contained in a finite subring. A topological  ring $(R,\mathcal T)$  is called {\it metrizable} if its underlying additive group satisfies the first axiom of countability. A ring $R$ with $1$ is called {\it Dedekind-finite} if each equality $xy=1$ implies $yx=1$. It is well-known that every finite ring with identity is Dedekind-finite. Since every compact ring with
identity is a subdirect product of finite rings, it follows that every compact ring with identity is Dedekind-finite.
If $A\subseteq  R$, then  ${\rm Ann}_l(A):=\{x\in R\mid xA=0\}$. If $X,Y$ are the subsets of $R$, then $X\cdot Y:=\{xy\mid x\in X, y\in Y\}$.  A topological ring $R$ is called {\it compactly generated} (see \cite{Ursul}, Chapter II) if there exists a compact subset $K$ such that $R=\langle K\rangle$. If $(R,\mathcal T)$ is a topological ring and $I$ is an ideal of $R$, then the quotient topology of the factor ring $R/I$ is denoted by $\mathcal T/I$. If $K$ is a  subset of  an abelian group $A$, then  set
\[
T(K)=\{\alpha\in {\rm End}\,(A)\mid \alpha(K)=0 \}.
\]
When $K$ runs over all finite subsets of $A$, the family $\{T(K)\}$ defines a ring topology ${\mathcal{T}_{fin}}$ on
${\rm End}\,(A)$. This topology is called the {\it finite topology}.

\begin{lemma} \label{LM:1}
For any abelian group $A$ the ring $({\rm End }\, A,\mathcal T_{fin})$ is complete.
\end{lemma}
\begin{proof}
See \cite{Ursul}, Theorem 19.2.
\end{proof}

\begin{lemma}\label{LM:2}
(Cauchy's criterion) In a Hausdorff complete commutative group $G$, in order that a family $(x_\alpha)_{\alpha\in\Omega}$ should be summable it is necessary and sufficient that, for each neighborhood $V$ of zero in $G$, there is a finite subset $\Omega_0$ of $\Omega$ such that $\Sigma_{\alpha\in K}x_\alpha\in V$ for all finite subsets $K$ of $\Omega$ which do not meet $\Omega$.
\end{lemma}
\begin{proof}
See \cite{Bourbaki_III}, p.263.
\end{proof}

\begin{lemma}\label{LM:3}
If $(x_\alpha)_{\alpha\in\Omega}$ is a summable subset in $({\rm End }\, A,\mathcal T_{fin})$ then  every subset $\Delta$ of $\Omega$ the family $(x_\beta)_{\beta\in\Delta}$ is summable.
\end{lemma}
\begin{proof}
Let $V$ be a neighborhood of zero of $({\rm End }\, A,\mathcal T_{fin})$. We can consider without loss of generality that $V$ is a left ideal of ${\rm End }\, A$. There exists a finite subset $\Omega_0$ of $\Omega$ such that $\Sigma_{\alpha\in K}x_\alpha\in V$ for every finite subset $K$ of $\Omega$ for which $K\cap\Omega_0=\varnothing$. Let $F$ be a finite subset of $\Delta$ such that $F\cap(\Omega_0\cap \Delta)=\varnothing$. If $\alpha\in F$, then $\alpha\notin\Omega_0$, hence $\Sigma_{\alpha\in F}x_\alpha\in V$. By Cauchy's criterion the family $(x_\beta)_{\beta\in\Delta}$ is summable.
\end{proof}

A topological ring $(R,\mathcal{T})$ is called {\it minimal} (see, for instance, \cite{Dikranjan})  if there is no ring topology $\mathcal{U}$ such that $\mathcal{U}\leq \mathcal{T}$
and $\mathcal{U}\not= \mathcal{T}$. A topological ring $(R,\mathcal{T})$ is called {\it simple} if  $R$ is simple as a ring without topology. A topological ring $(R,\mathcal{T})$ is called {\it weakly simple}  if  $R^2\neq 0$ and  every its closed ideal is either  $0$ or $R$.
A topological ring $(R,\mathcal{T})$ is called {\it completely simple} (see  \cite{Mutylin})  if  $R^2\neq 0$ and  for every continuous homomorphism   $f: (R,\mathcal{T})\to(S,\mathcal{U})$ in a topological ring $(S,\mathcal U)$ either  $\ker(f)=R$ or $f$ is a homeomorphism of $(R,\mathcal T)$ on ${\rm Im}(f)$. Equivalently,   $R^2\not= 0$ and   $(R,\mathcal{T})$ is weakly simple and minimal.
Let $M$ be a unitary right $R$-module over   a commutative ring $R$  with $1$.
The  module $M$ is called \emph{divisible} if $Mr=M$ for every $0\neq r\in R$.
A  right  $R$-module $M$ is called {\it faithful} if $Mr=0$ implies $r=0$ ($r\in R$).
A   right  $R$-module $M$ is called {\it torsion-free} if $mr=0$ implies that either $m=0$ or $r=0$, where $m\in M$ and $r\in R$.
Recall that a submodule $N$ of an $R$-module $M$ is called {\it fully invariant} $\alpha(N)\subseteq N$ for every endomorphism $\alpha$ of $M_R$.
We use in the sequel the notion and results from  the books \cite{Engelking, Ursul}.

\begin{remark}
If $R$ is a von Neumann regular ring, then $R^2=R$.
\end{remark}

\begin{lemma}\label{LL:1}
An ideal $I$ of a von Neumann regular ring is von Neumann regular.
\end{lemma}

\begin{proof}
Let $i\in I$. Thus there exists $x\in R$ such that $ixi=i$. It follows that $ixixi=i $ and $xix\in I$.
\end{proof}

\begin{corollary}If $I$ an ideal of  a von Neumann regular ring $R$, then any ideal $H$ of $I$ is an ideal of $R$, too.
\end{corollary}
\begin{proof}
$RH=RH^2\subseteq IH\subseteq H$. Similarly, $HR\subseteq H$.
\end{proof}

If $A_p$ is a $p$-elementary countable group, then
\[
I=\{\alpha\in {\rm End}\,(A_p)\vert\vert  {\rm Im}(\alpha)\vert<\aleph_0\}.
\]
Fix a linear basis $\{v_i\mid i\in \mathbb N\}$ of $A_p$ over the field $\mathbb F_p$. Using this fixed  basis, we define  the map  $e_i: A \to A$ such that
\[
e_i(v_j)=\delta_{ij}v_j, \qquad (i, j\in\mathbb N)
\]
where $\delta_{ij}$ is the Kronecker delta.

\begin{lemma}\label{LL:2} We have for ${\rm End}\,(A_p)$:
\begin{itemize}
\item[(i)] $I$ is a von Neumann regular ring.
\item[(ii)] $I$ is a simple ring.
\item[(iii)] The factor ring ${\rm End}(A_p)/I$ is  simple  von Neumann regular.
\item[(iv)] $I$ is a locally finite ring.
\end{itemize}
\end{lemma}
\begin{proof}

(i):\quad  It is well-known that  ${\rm End}(A_p)$ is regular (see \cite{Lam}, Theorem 4.27, p.63). Then  $I$ is von Neumann regular by Lemma \ref{LL:1}.

(ii), (iii):\quad  The ideal  $I$ is the only nontrivial ideal of   ${\rm End}(A_p)$ (see \cite{Jacobson}, \S17, Theorem 1,  p.93).  This means that  ${\rm End}(A)/I$  is simple. It is  regular by the part (i).

(iv)\quad  Since $I$ is simple (see \cite{Jacobson},  \S 12, Proposition 1),  it suffices to show that $I$ contains a nonzero locally finite right ideal.

Let us show  that the left ideal $Ie_1$ of $I$ is locally finite as a ring  (equivalently, as a $\mathbb F_p$-algebra). We have $0\neq e_1\in Ie_1$.  If $H$ is the left annihilator of $Ie_1$, then, obviously, $H$ is a locally finite ring, hence it is locally finite as a $\mathbb F_p$-algebra. We claim that $Ie_1/H$ is finite.
Indeed, define $\beta_n\in H$ $(n\ge 2)$ in the following way
\[
\beta_n(v_i)=
\begin{cases} v_n,\quad\text{for}\quad i=1;\\
0,\qquad  \text{for}\quad i\neq 1.\\
\end{cases}
\]
We claim that $Ie_1=\mathbb F_pe_1+\Sigma_{n=2}^\infty \mathbb F_p\beta_n$.

For, let $\alpha\in I$. Then $\alpha(v_1)=r_1v_1+\cdots+r_nv_n$, where $r_i\in\mathbb F_p$ and $n\in\mathbb N$.  Thus
\[
\begin{split}
\alpha e_1(v_1)& =r_1e_1(v_1)+r_2\beta_2(v_1)+\cdots+r_n\beta_n(v_1)\\
&=(r_1e_1+r_2\beta_2+\cdots+r_n\beta_n)(v_1);\\
\alpha e_1(v_j)&=(r_1e_1+r_2\beta_2+\cdots+r_n\beta_n)(v_j)\qquad (j\neq 1).
\end{split}
\]
This yields
\[
\alpha e_1=r_1e_1+r_2\beta_2+\cdots+r_n\beta_n \qquad
\]
and so $Ie_1=\mathbb F_pe_1+\Sigma_{n=2}^\infty \mathbb F_p\beta_n$.

In particular,
$Ie_1=\mathbb F_pe_1+H$, and so $H$ has a finite index in $Ie_1$. Clearly,  $Ie_1$ is a locally finite $\mathbb F_p$-algebra (see \cite{Jacobson},  Proposition 1,  p.241) and  $I$ is a locally finite $\mathbb F_p$-algebra (see \cite{Jacobson}, Proposition 2, p.242).\end{proof}

The next result  can be deduced from Lemma 36.11 of \cite{Ursul}.
\begin{lemma}\label{LL:3}
Let $A$ be a  locally compact, compactly generated, and  totally disconnected ring. If  $A$ contains a dense   locally finite subring $B$, then $A$ is compact.
\end{lemma}
\begin{proof}
Let $A=\langle V\rangle$, where $V$ is a compact symmetric neighborhood of zero. Since $V$ is compact, the subset $V+V+V\cdot V$ also is compact. Since $B$ is dense, $A=B+V$. By compactness of $V+V+V\cdot V$ there exists a 	 finite subset $H\subseteq  B$ such that $V+V+V\cdot V\subseteq H+V$. Since $B$ is a locally finite ring, we can assume without loss of generality that $H$ is a subring. Let  $H\setminus\{0\}=\{h_1,\ldots,h_k\}$.
The subset
\[
H+h_1V+\cdots+h_kV+V
\]
 is an open subgroup of $R(+)$. Indeed, this subset is symmetric and
\[
\begin{split}
(H+h_1V+\cdots+h_kV+V)+(H+h_1V+\cdots+h_kV+V)&\\
\subseteq H+h_1(V+V)+\cdots+h_k(V+V)+V+V&\\
\subseteq H+h_1V+\cdots+h_kV+V&.
\end{split}
\]
We  prove by induction on $m$ that
\[
V^{[m]}\subseteq H+h_1V+\cdots+h_kV+V,  \qquad (m\in\mathbb{N})
\]
where $V^{[1]}=V$ and $V^{[m]}=V^{[m-1]}\cdot V$ for all $m$.

The inclusion is obvious for $m=1$.

Assume that the assertion has been proved for $m\ge1$. Clearly,
\[
\begin{split}
V^{[m+1]}=V^{[m]}\cdot V\subseteq H\cdot V+h_1(V\cdot V)+\cdots + h_k(V\cdot V)+V\cdot V& \subseteq \\
h_1V+\cdots+ h_kV+h_1(H+V)+\cdots+ h_k(H+V)+H+V& \subseteq\\
H+h_1V+\cdots+h_kV+V&.
\end{split}
\]
Consequently,   $A=H+h_1V+\cdots+h_kV+V$, therefore $A$ is compact.
\end{proof}

An element $x$ of a topological ring is called {\it discrete} if there exists a neighborhood $V$ of zero such that $xV=0$ (i.e., the right annihilator of $x$ is open).

\begin{lemma}\label{LL:4}
The set of all discrete elements of  a topological ring is  an ideal. A simple ring with identity does not contain nonzero discrete elements.
\end{lemma}

\section{Locally compact ring topologies  on  ${\rm End}\,(A)$ of a countable elementary abelian $p$-group $A$}\label{P:1}

\begin{theorem}\label{T:1}
Let $R$ be a simple,   nondiscrete and  locally compact ring of  ${\rm char}(R)=p>0$ and $1\in R$.
If $V$ is a compact open subring of $R$ and $\{e_\alpha\vert \alpha\in\Omega\}$ is a set of orthogonal idempotents in $R$, then
\[
|\Omega|\leq w(V).
\]
\end{theorem}
\begin{proof}
The ring $R$ does not contain nonzero discrete elements by Lemma  \ref{LL:4}. Since $R$ is locally compact and ${\rm char}(R)=p$,  it is totally disconnected. Additionally, $R$ has a fundamental system of neighborhoods of zero consisting of compact open subrings by  Lemma 9 of \cite{Kaplansky}.

If $V$ is a compact open subring of $R$, then by continuity of the ring operations  for each $\alpha\in\Omega$ there exists an open ideal $V_\alpha$ of $V$ such that $e_\alpha V_\alpha\subseteq V$.  Clearly, there exists   $y_\alpha\in V_\alpha$ for which   $e_\alpha y_\alpha\neq 0$ since $R$ has no nonzero discrete elements.

\noindent
We claim that hold the following two properties:

(i) $e_\alpha y_\alpha\not\in \overline{\{e_\beta y_\beta\vert \beta\neq \alpha\}}$ for each $\alpha\in\Omega$;

(ii) the set    $X=\{e_\alpha y_\alpha\vert \alpha\in \Omega\}$ is a discrete subspace of $V$.

Indeed, if were  $e_\alpha y_\alpha\in \overline{\{e_\beta y_\beta\vert \beta\neq \alpha\}}$ for some $\alpha\in \Omega$, then
\[
\begin{split}
e_\alpha y_\alpha=e_\alpha e_\alpha y_\alpha&\in e_\alpha\overline{\{e_\beta y_\beta\vert \beta\neq \alpha\}}\\
&\subseteq \overline{\{e_\alpha e_\beta y_\beta\vert \beta\neq \alpha\}}\\
&=\{0\},
\end{split}
\]
so  $e_\alpha y_\alpha=0$, a contradiction. The part (i) is proved.

(ii) Now,   for  each $\alpha\in\Omega$  we have  $V\setminus\overline{\{e_\beta y_\beta\vert \beta\neq \alpha\}}$ is open and, consequently,
\[
(V\setminus\overline{\{e_\beta y_\beta\vert \beta\neq \alpha\}})\cap X=\{e_\alpha y_\alpha\},
\]
by (i). Therefore the point $e_\alpha y_\alpha(\alpha\in\Omega)$ of $X$ is isolated. In other words, the subspace $X$ of $V$ is discrete.

Since $X$ is discrete,      $|\Omega|=|X|=w(X)\leq w(V)$  (see \cite{Arkhangelskii_Ponomarev}, Exercises 98-99, p.72).
\end{proof}

\begin{theorem}\label{T:2}
Let $A_p$ be a countable elementary abelian $p$-group.
Then the ring \[
I=\{\alpha\in {\rm End}\,(A_p)\vert\vert  {\rm Im}(\alpha)\vert<\aleph_0\}
\]
 does not admit a nondiscrete  ring topology $\mathcal U$
such that $(I,\mathcal U)$ is a Baire space.
\end{theorem}
\begin{proof}

Put $S_n=\{\alpha\in I\mid  \alpha(A)\subseteq {\mathbb{F}_p} v_1+\cdots+\mathbb F_pv_n\}$, where $n\in\mathbb{N}$.
Clearly  $I=\cup_{n\in\mathbb{N}}S_n$ and
\[
S_n=\{\alpha\in I \mid e_i\alpha=0 \quad  \text{for} \quad  i> n\}={\rm Ann}_r\big(\{e_k\mid k>n\}\big).
\]
This yields that   $S_n$ is closed due continuity of the ring operations.

Since $I$ is a Baire space, there exists $n\in\mathbb{N}$ such that ${\rm Int}(S_n)\neq\emptyset$, hence $S_n$ is an open subgroup.

Set $\beta\in I$ such that
\[
\beta(v_i)=
\begin{cases} v_{n+i},\quad\text{for}\quad i=1,\ldots,n;\\
0\quad,  \quad \text{for}\quad i> n.\\
\end{cases}
\]
Let $W\subseteq S_n$ be a neighborhood of zero of $(I,\mathcal U)$ such that $\beta W\subseteq S_n$. If $w\in W\setminus 0$,  then there exist $a\in A$ and $ r_1,\ldots, r_n\in\mathbb F_p$ such that
\[
0\neq w(a)=\sum_{i=1}^n r_iv_i\qquad
\text{and} \qquad \beta( w(a))=\sum_{i=1}^n  r_iv_{n+i}.
\]
There exists  $j\in 1,\ldots, n$ such that $r_j\neq 0$. Then
\[
e_{n+j}\beta w(a)=r_jv_{n+j}\neq 0,
\]
 hence $e_{n+j}\beta w\neq 0$ and so  $\beta w\not\in S_n$, a contradiction.\end{proof}

\begin{corollary}\label{C:1}
Under the notation of Theorem \ref{T:2} the ring  $I$ does not admit a nondiscrete locally compact ring topology.
\end{corollary}
\begin{proof}
This follows from the fact that  each  locally compact space is a Baire space  (see \cite{Bourbaki}, Theorem 1, p.117).
\end{proof}

Our main result is the following.

\begin{theorem}\label{T:3}
The endomorphism ring ${\rm End}\,(A_p)$ of a countable elementary abelian $p$-group $A_p$ does not admit a nondiscrete locally compact ring topology.
\end{theorem}
\begin{proof}

We will use the notation and results from section 2.
Denote $R= {\rm End}\,(A_p)$. Assume on the contrary that there exists on $R$ a nondiscrete locally compact ring topology $\mathcal T$.

\noindent
\underline{Fact 1.} The ring $(R,\mathcal T)$ has a fundamental system of neighborhoods of zero consisting of compact open subrings.

Since  the additive group of $R$  has exponent $p$, it  is totally
disconnected (this follows from \cite{Hewitt_Ross}, Theorem 9.14, p.95).
By   I.~Kaplansky's result (see \cite{Kaplansky}, Lemma 9) $(R,\mathcal T)$ has a  fundamental system of neighborhoods of zero consisting of compact open subrings.

\noindent
\underline{Fact 2.} The group  $Re_i$
is countable for each $i\in\mathbb N$.

\noindent
We claim  that $Re_i$ is infinite. Indeed, for each
$j\in \mathbb{N}$  put $\beta_j\in R$ such that
\[
\beta_j(v_k)=
\begin{cases}
v_j, \quad \text{for}\quad k=i;\\
0, \qquad \text{for}\quad  k\not=i.\\
\end{cases}
\]
If $j\not=s$, then $\beta_je_i(v_i)=\beta_j(v_i)=v_j$ and
$\beta_se_i(v_i)=\beta_s(v_i)=v_s$, hence $\beta_je_i\not=\beta_se_i$, so $Re_i$
is infinite.

The ring $Re_i$ is countable. Indeed, consider the mapping $\psi: Re_i\to A_p^{\mathbb{F}_pv_i}$, where
\[
\psi(\alpha e_i)(rv_i)=\alpha(rv_i)\qquad\text{for all}\qquad  r\in \mathbb{F}_p.
\]
If $\alpha e_i\not=\beta e_i$ ($\alpha, \beta\in R$), then there exists $x=\sum_j r_jv_j\in A_p$ such that
$\alpha e_i(x)\not=\beta e_i(x)$, hence,  $\alpha (r_iv_i)\not=\beta(r_iv_i)$. Thus
\[
\psi(\alpha e_i)(r_iv_i)=\alpha (r_iv_i)\not=\beta(r_iv_i)=\psi(\beta e_i)(r_iv_i).
\]
The latter means that $\psi$ is an injective mapping of  $Re_i$ into $A^{\mathbb{F}_pv_i}$ .
Since $A^{\mathbb{F}_pv_i}$ is countable, $Re_i$ is countable, too.

\noindent
\underline{Fact 3}. $I$ is a closed ideal of $R$.
 We claim that $I$ is not dense in  $(R,\mathcal T)$.
Assume the contrary.
Since $I$  is locally  finite and is a maximal ideal,   $(R,\mathcal T)$
is topologically locally finite by Lemma  \ref{LL:3}. The ring $R$ contains two elements
$x, y$ such that $xy=1$ and $yx\neq 1$. The subring $\overline{\langle x,y\rangle}$
is compact, hence Dedekind-finite, a contradiction. We obtained that
$(R/I,\mathcal T/I)$ is a nondiscrete  metrizable locally compact ring.

\noindent
\underline{Fact 4.} $I$ is a discrete ideal of $R$.

\noindent
This follows from Theorem 2.

\noindent
\underline{Fact  5.} $Re_i$ is a discrete left ideal of $R$ for every $i\in\mathbb N$.

\noindent
Indeed, $Re_i\subseteq I$ and $I$ is discrete by Fact 4 for every $i\in\mathbb N$.

\noindent
\underline{Fact  6.}  ${\rm Ann}_l(e_i)$ is  open in $R$ for every $i\in\mathbb N$.

\noindent
 Indeed, the group homomorphism
$q: R\to Re_i, r\mapsto re_i,$ is continuous. Since $Re_i$ is discrete $q^{-1}(0)= {\rm Ann}_l(e_i)$
is open.

\noindent
\underline{Fact 7.}   $\cap_i{\rm Ann}_l(e_i)=0$.

\noindent
Obvious.

\noindent
\underline{Fact 8.} $\mathcal{T}\geq \mathcal{T}_{fin}$.

We notice that ${\rm Ann}_l(e_i)=T(\{v_i\})$ for every $i\in\mathbb N$. For, if $\alpha e_i=0$,  then $\alpha(v_i)=\alpha e_i(v_i)=0$, i.e., $\alpha\in T(\{v_i\})$.
Conversely, if  $\alpha\in T(\{v_i\})$, then $\alpha e_i(v_i)=\alpha(v_i)=0$. If $j\not=i$ then $\alpha e_i(v_j)=0$. Therefore
$\alpha e_i=0$. Furthermore
\[
T(\{v_1,\ldots,v_n\})=\cap_{i=1}^n T(\{v_i\})=\cap_{i=1}^n {\rm Ann}_l(e_i)\in \mathcal{T}
\] for each $n\in\mathbb N.$\\
Since the family  $\{T(\{v_1,\ldots,v_n\})\}$ forms a fundamental system of  neighborhoods
of zero of $(R,\mathcal T_{fin})$, we get that $\mathcal T_{fin}\leq \mathcal{T}$.

\noindent
\underline{Fact 9.} The ring $(R,\mathcal T)$  is metrizable.

Since $\cap_{i\in\mathbb{N}}{\rm Ann}_l(e_i)=0$, the pseudocharacter of $(R,\mathcal T)$ is $\aleph_0$.
If $V$ is a compact open subring of $(R,\mathcal T)$ (see Fact 1), then the pseudocharacter of $V$ also is $\aleph_0$.
However in every compact space the pseudocharacter of a point coincides with its character. Therefore $(R,\mathcal T)$  is metrizable.

\noindent
\underline{Fact 10.}
$(R/I,\mathcal T/I)$ has an open compact subring.

Indeed, it is well-known (see \cite{Kaplansky}) that every totally disconnected ring has a fundamental system of
neighborhood of zero consisting of compact open subrings.
Henceforth  $V$ is  a fixed open compact  subring of $(R/I,\mathcal T/I)$.

\noindent
\underline{Fact 11.}  $R/I$ contains a family of  orthogonal idempotents of cardinality $2^{\aleph_0}$ .

Indeed, the family $\{e_i\}_{i\in\mathbb{N}}$ of idempotents of  $(R, \mathcal T_{fin})$ is summable in  and $1_A=\Sigma_{n\in\mathbb{N}}e_n$, where $1_A$ is the identity of $R$.

The first ordinal  number of cardinality $\mathfrak{c}$ of continuum is denoted by $\omega(\mathfrak{c})$. Let $\{\mathbb{N}(\alpha)\vert \alpha< \omega(\mathfrak{c})\}$  be a family of infinite almost disjoint subsets of $\mathbb{N}$ (see \cite{Engelking}, Example 3.6.18, p.175--176).
Put  $f_{\mathbb{N}(\alpha)}=\Sigma_{i\in \mathbb{N}(\alpha)}e_i$ for every $\alpha<\omega(\mathfrak{c})$. The element $f_{\mathbb{N}(\alpha)}$ exists by Lemma \ref{LM:3}. Then:
\begin{itemize}
\item[(i)] $f_{\mathbb{N}(\alpha)}\notin I$ for every $\alpha  <\omega(\mathfrak{c})$;
\item[(ii)] $f_{\mathbb{N}(\alpha)}f_{\mathbb{N}(\beta)}\in I$ for each $\alpha,\beta <\omega(\mathfrak{c})$ and $ \alpha\neq \beta$.
\end{itemize}
If  $g_\alpha=f_{\mathbb{N}(\alpha)}+I$ for each $\alpha< \omega(\mathfrak{c})$,  then $\{g_\alpha\vert \alpha <\omega(\mathfrak{c})\}$ is the required system of orthogonal idempotents.

The subring  $V$ is metrizable (by Fact 9). Since $V$ is  compact and $R/I$ is a simple von Neumann regular  ring by Lemma  \ref{LL:2} and  $w(V)\leq \aleph_0$,  we obtain a   contradiction to Theorem \ref{T:1}. \end{proof}

\begin{theorem}\label{T:4}(CH)
Under the notation of Theorem \ref{T:3},  the ring $R/I$ does not admit a nondiscrete locally compact ring topology.
\end{theorem}
\begin{proof}
Assume on the contrary that  the factor ring $R/I$ admits  a nondiscrete locally compact ring topology $\mathcal{T}$, so $(R/I, \mathcal{T})$ contains an open compact subring $V$. Since the cardinality of $R/I$ is continuum and $V$ is infinite, the power of $V$ is continuum. Since we have assumed (CH), the subring  $V$ is  metrizable, hence second metrizable (see \cite{Hulanicki, Jones}).  However  we have proved in  Theorem \ref{T:3} that the ring  $R/I$ contains a family of orthogonal idempotents of cardinality $\mathfrak{c}$, a  contradiction with Theorem \ref{T:1}.\end{proof}

\begin{theorem}\label{T:5}
The finite topology $\mathcal T_{fin}$ is the only second metrizable ring topology $\mathcal T$ on $R$  for which $(R,\mathcal T_{fin})$ is complete.
\end{theorem}
\begin{proof}
Let $K=\langle F\rangle$, where  $F$ is  a finite subset of $A$. Clearly,  there exists a subgroup $A^\prime$ of $A$ such that $A=K\oplus A^\prime$.
Choose  ${\rm e}_F\in R$ such that ${\rm e}_F\upharpoonright_K={\rm id}_K$ and ${\rm e}_F(A^\prime)=0$.  Clearly
\[
T(K)=R(1-e_F)
\]
and   $\alpha K=0$ if and only if  $\alpha\in R(1-e_F)$. It follows that the family
$\{R(1-e_F)\}$ where  $F$ runs over all finite subset of $A$, forms a fundamental system
of neighborhoods of zero for $(R,\mathcal T_{fin})$.

There exists an injective map of $Re_F$ to ${\rm Hom}(K,A)$, so the left ideal $Re_F$
is countable, due to countability  ${\rm Hom}(K,A)$. Since $e_F^2=e_F$, the Peirce decomposition
\[
R=Re_F\oplus R(1-e_F)
\]
of $R$ with respect to the idempotent $e_F$ is a decomposition of the  topological group $(R,+,\mathcal T)$.  It follows that $Re_F$ is discrete, hence $R(1-e_F)$ is open (in the topology $\mathcal T$). It follows that $\mathcal T\ge\mathcal T_{fin}$, so $\mathcal T=\mathcal T_{fin}$ (see \cite{Freudenthal},  Theorem 30    or \cite{Graev}). \end{proof}

\section{Completely simple topological endomorphism rings\\ of vector spaces}\label{P:2}

\begin{theorem}\label{T:6}
Let $A_F$ be a right vector space over a division ring $F$ and $S={\rm End}\,(A_F)$.
The following conditions are equivalent:
\begin{itemize}
\item[\rm(i)] $(S, \mathcal{T}_{fin})$ is a completely simple
topological ring.
\item[\rm(ii)] $\dim(A_F)=\infty$ or $\dim(A_F)< \infty$ and $F$ does not admit a nondiscrete ring topology.
\end{itemize}
\end{theorem}
\begin{proof}
(i)$\Rightarrow$ (ii): If $A_F$ is finite-dimensional, then $S$ is discrete and isomorphic to
$\rm M(n,F)$, where $n$ is the dimension of $A_F$. Then, obviously, $F$ does not admit a
nondiscrete ring topology.

(ii)$\Rightarrow$ (i):
If $\dim(A_F)= n< \infty$, then  $S\cong \rm M(n,F)$. Since $F$ does not admit nondiscrete ring topologies, the same holds for  $\rm M(n,F)$.

Let $A_F$ be  infinite dimensional. Fix a   basis $\{x_\alpha\}_{\alpha<\tau}$ over $F$, where $\tau$
is an infinite ordinal number. It is well-known  that the topological ring   $(S, \mathcal{T}_{fin})$ is weakly  simple (see \cite{Leptin}, Satz 12, p.258) and  the family $\{T(x_\alpha)\}_{\alpha<\tau}$ is a prebase at
zero for the finite topology $\mathcal T_{fin}$ of $S$.

Assume on the contrary that there exists a Hausdorff ring topology $\mathcal T$,
coarser that $\mathcal T_{fin}$ and different from  it. Let $e_\alpha\in S$ such that $e_\alpha^2=e_\alpha$ and  $e_\alpha(x_\beta)=\delta_{\alpha\beta}x_\alpha$ for each $\alpha<\tau$, where  $\delta_{\alpha\beta}$ is the Kronecker delta.

\noindent
\underline{Fact 1}. $T(x_\alpha)={\rm Ann}_l(e_\alpha)$ for each $\alpha<\tau$.

Indeed, if $p\in T(x_\alpha)$, then $pe_\alpha(x_\alpha)=p(x_\alpha)=0$. If $\beta \neq \alpha$, then $e_\alpha(x_{\beta})=0$, hence $pe_\alpha=0$, i.e. $p\in {\rm Ann}_l(e_\alpha)$.
Conversely, if $pe_\alpha=0$, then $p(x_\alpha)=pe_\alpha(x_\alpha)=0$, i.e.  $p\in T(x_\alpha)$.

\noindent
\underline{Fact 2}. There exists $\alpha_0<\tau$ for which $Se_{\alpha_0}$ is nondiscrete in $(S,\mathcal T)$.

Assume on the contrary that   for every $\alpha<\tau$ there exists a neighborhood $V_\alpha$ of zero of $(S,\mathcal T)$ such that $Se_\alpha\cap V_\alpha=0$. If $U_\alpha$ is a neighborhood of zero of  $(S,\mathcal T)$ such that $U_\alpha e_\alpha\subseteq V_\alpha$, then $U_\alpha e_\alpha=0$, hence ${\rm Ann}_l(e_\alpha)=T(x_\alpha)$ is open in $(S,\mathcal T)$. Hence  $\mathcal{T}_{fin}\leq\mathcal T$ and  $\mathcal T=\mathcal T_{fin}$, a contradiction.

\noindent
\underline{Fact 3}. \quad $(Se_{\alpha_0}\cap V)x_{\alpha_0}\nsubseteq \oplus_{\beta\in K}x_\beta F$

\noindent
for any neighborhood $V$ of of zero of $(S,\mathcal T)$ and any finite subset $K$ of the set $[0,\tau)$ of all ordinal numbers less than $\tau$.

Assume on the contrary that there exists a finite subset $K$ of  $[0,\tau)$ and a neighborhood $V$ of zero of $(S,\mathcal T)$ such that
\begin{equation}\label{E:1}
(Se_{\alpha_0}\cap V)x_{\alpha_0}\subseteq \oplus_{\beta\in K}x_\beta F.
\end{equation}
Fix $\gamma\in [0,\tau)\setminus K$. For each $\beta\in K$ define  $q_\beta\in S$ such that $q_\beta(x_\beta)=x_\gamma$ and $q(x_\delta)=0$ for $\delta\not=\beta$.

Let $V_0$ be a neighborhood of zero of $(S,\mathcal T)$ such that $V_0\subseteq V$ and $q_\beta V_0\subseteq V$ for all $\beta\in K$.
There exists $0\neq h\in Se_{\alpha_0}\cap V_0$ by  Fact 2. It follows that $hx_{\alpha_0}\not=0$ by  Fact 1.
Since $Se_{\alpha_0}\cap V_0\subseteq Se_{\alpha_0}\cap V$, we obtain that  $hx_{\alpha_0}=\Sigma_{\beta\in K}x_\beta f_\beta$, ($f_\beta\in F$)  by (\ref{E:1}). There exists $\beta_0\in K$ such that $f_{\beta_0}\neq 0$ (because $hx_{\alpha_0}\not=0$), so
\[
\textstyle
q_{\beta_0}h=q_{\beta_0}(\Sigma_{\beta\in K}x_\beta f_\beta)=r_{\beta_0}x_\gamma\not\in \oplus_{\beta\in K}x_\beta F,
\]
a contradiction. Therefore  Fact 3 is proved.

Now let $V$ be a  neighborhood of zero of $(S,\mathcal T)$. Pick up    a neighborhood $V_0$ of zero of $(S,\mathcal T)$  such that $V_0\cdot V_0\subseteq V$.
Since $\mathcal{T}\leq\mathcal{T}_{fin}$,   there exists a finite subset $K$ of $[0,\tau)$ such that
\[
T(\{x_\beta\mid \beta\in K\})\subseteq V_0.
\]
We have   $(Se_{\alpha_0}\cap V_0)x_{\alpha_0}\nsubseteq \oplus_{\beta\in K}x_\beta F$ by  Fact 3. It follows that there exists  $q\in Se_{\alpha_0}\cap V_0$ such that
\[
q(x_{\alpha_0})\not\in \oplus_{\beta\in K}x_\beta F.
\]
Clearly,  $q(x_{\alpha_0})\in A_F$, so it  can be written as  $q(x_{\alpha_0})=\sum_{\alpha< \tau}x_{\alpha}f_{\alpha}$, where $f_{\alpha}\in F$ and here exists $\beta_0\not\in K$  such that $f_{\beta_0}\not=0$.

Consider the element   $s\in S$ such that  $s(x_{\beta_0})=x_{\alpha_0}f_{\beta_0}^{-1}$ and $s(x_{\lambda})=0$ for $\lambda\neq\beta_0$. Evidently, $s\in T(K)$, hence
\[
sq\in T(K)\cdot V_0\subseteq V_0\cdot  V_0\subseteq V.
\]
Moreover, $sq(x_{\alpha_0})=s(x_{\beta_0}f_{\beta_0}+\cdots)=x_{\alpha_0}$. Since $q\in Se_{\alpha_0}$ we obtain that $sq(x_\beta)=0$ for $\beta\neq \alpha_0$. Consequently,  $e_{\alpha_0}=sq\in V$ for every neighborhood $V$ of zero of $(S,\mathcal T)$, a contradiction.
\end{proof}
\begin{remark}
The question of existence of a uncountable division ring which does not admit a nondiscrete Hausdorff ring topology is open.
Several results on this topic  can be found in Chapter 5 of \cite{Arnautov_book}.
\end{remark}

\begin{theorem}\label{T:7}
Let  $\prod_{\alpha\in \Omega}R_{\alpha}$ be a family of compact rings with identity. Then the product $(\prod_{\alpha\in\Omega}R_\alpha,\prod_{\alpha\in\Omega}\mathcal T_\alpha)$ is a minimal ring if and only if  every $(R_\alpha,\mathcal T_\alpha)$ is a minimal topological ring. (Here $\prod_{\alpha\in\Omega}\mathcal T_\alpha$ is the product topology on $\prod_{\alpha\in\Omega}R_\alpha$.)
\end{theorem}
\begin{proof}
$\Rightarrow:$ Assume on the contrary that there exists $\beta\in\Omega$ and a ring topology $\mathcal T^\prime$ on $R_\beta$ such that $\mathcal T^\prime\leq\mathcal T_\beta$ and $ \mathcal T^\prime\neq\mathcal T_\beta$. Consider the product topology $\mathcal U$ on $\prod_{\alpha\in\Omega}R_\alpha$, where $R_\alpha$ is endowed with $\mathcal T_\alpha$ when $\alpha\neq \beta$ and $R_\beta$ is endowed with $\mathcal T^\prime$. Obviously, $\mathcal U\leq\prod_{\alpha\in\Omega}\mathcal T_\alpha$ and $\mathcal U\neq\prod_{\alpha\in\Omega}\mathcal T_\alpha$, a contradiction.

\noindent
$\Leftarrow:$
Denote by $\pi_\alpha(\alpha\in\Omega)$ the projection of $\prod_{\alpha\in \Omega}R_{\alpha}$ on $R_\alpha$.  By definition of the product topology, $\prod_{\alpha\in \Omega}\mathcal T_{\alpha}$ is the coarsest topology on $\prod_{\alpha\in \Omega}R_{\alpha}$ for which the projections $\pi_\alpha(\alpha\in\Omega)$ are continuous.

Let $\mathcal U$  be a ring topology on  $\prod_{\alpha\in \Omega}R_{\alpha}$, $\mathcal U\leq\prod_{\alpha\in \Omega}\mathcal T_\alpha$ and $\beta\in \Omega$. Since
\[
\mathcal U\upharpoonright_{R_\beta\times\prod_{\gamma\neq\beta}\{0_\gamma\}}\leq\big(\prod_{\alpha\in\Omega}\mathcal T_\alpha\big)\upharpoonright_{R_\beta\times\prod_{\gamma\neq\beta}\{0_\gamma\}},
\]
it follows that $\mathcal U\upharpoonright_{R_\beta\times\prod_{\gamma\neq\beta}\{0_\gamma\}}=(\prod_{\alpha\in\Omega}\mathcal T_\alpha)\upharpoonright_{R_\beta\times\prod_{\gamma\neq\beta}\{0_\gamma\}}$ by minimality of $(R_\beta,\mathcal T_\beta)$.

Then the family $\{V\times\prod_{\gamma\neq\beta}\{0_\gamma\}\}$ when $V$ runs all neighborhoods of zero of $(R_\beta,\mathcal T_\beta)$ is  a fundamental system of neighborhoods of zero of
\[
\Big(R_\beta\times\prod_{\gamma\neq\beta}\{0_\gamma\},\quad \mathcal U\upharpoonright_{R_\beta\times\prod_{\gamma\neq\beta}\{0_\gamma\}}\Big).
\]
Since $R_\beta\times\prod_{\gamma\neq\beta}\{0_\gamma\}$ is an ideal with identity of $\prod_{\alpha\in \Omega}R_{\alpha}$, the topological ring $(\prod_{\alpha\in \Omega}R_{\alpha},\mathcal U)$ is a direct sum of ideals $R_\beta\times\prod_{\gamma\neq\beta}\{0_\gamma\}$ and $\{0_\beta\}\times\prod_{\gamma\neq\beta}R_\gamma$.
Let  $V$ be a neighborhood of zero of $(R_\beta,\mathcal T_\beta)$.
Then $V\times \prod_{\gamma\neq\beta}R_\gamma$ will be a neighborhood of zero of $(\prod_{\alpha\in \Omega}R_{\alpha},\mathcal U)$ and $\pi_\beta(V\times \prod_{\gamma\neq\beta}R_\gamma)=V$.

We have proved that $\pi_\beta$ is a continuous function from $(\prod_{\alpha\in \Omega}R_{\alpha},\mathcal U)$ to $(R_\beta,\mathcal T_\beta)$. It follows that  $\prod_{\alpha\in \Omega}\mathcal T_\alpha\leq \mathcal U$ and so $\mathcal U=\prod_{\alpha\in \Omega}\mathcal T_\alpha.$
\end{proof}

\begin{corollary}\label{C:2}
A left linearly compact semisimple ring is minimal if and only if has no direct summands of the form $M(n,\Delta)$, where $\Delta$ is a division ring which does not admit a nondiscrete Hausdorff ring topology.
\end{corollary}
\begin{proof}
This follows from Theorems \ref{T:6}, \ref{T:7} and the Theorem of Leptin (see \cite{Leptin}, Theorem 13, p. 258) about the structure of  left linearly compact semisimple rings.
\end{proof}

\begin{corollary}\label{C:3}
A semisimple linearly compact ring $(R, \mathcal{T})$ having no ideals isomorphic to matrix rings over infinite division rings is minimal.
\end{corollary}

%\newpage
\section{Completely simple endomorphism rings of modules}\label{P:3}

The  endomorphism ring of a right $R$-module $M$ is denoted by\\
 ${\rm End}\,(M_R)$.

\begin{lemma}\label{L:3}
Let $M$ be  a  divisible,  torsion-free module over a commutative domain $R$ and  $K$  the  field of fractions of $R$. The additive group of  $M$ has a structure of   a vector $K$-space such that $R$-endomorphisms of $M$ are exactly the  $K$-linear transformations.
\end{lemma}
\begin{proof}
We define a structure of a right vector $K$-space as follows: if  $\frac{a}{b}\in K$ and $m\in M$, then there exists a unique $x\in M$ such that $ma=xb$; set $m\circ \frac{a}{b} =x$.
Moreover,  if  $\frac{a}{b}=\frac{c}{d}$ and $0\neq m\in M$, then $m \circ \frac{a}{b}=m \circ \frac{c}{d}$. Indeed, if $m \circ \frac{a}{b}=x$ and $m \circ \frac{c}{d}=y$, then $mad=xbd$ and $mbc=ybd$ which means that $xbd=ybd$, hence $x=y$.

Let $\alpha\in {\rm End}\,(M_R)$, $\frac{a}{b}\in K$, $m\in M$. By definition, $am=b(\frac{a}{b}\circ m)$, hence, $a\alpha(m)= b\alpha(\frac{a}{b}\circ m)$, which means that $\alpha(\frac{a}{b}\circ m)=\frac{a}{b}\circ\alpha(m)$, so $\alpha$ is a $K$-linear transformation.
Note that,  if $a\in R$ and $m\in M$, then $m \circ \frac{a}{1}=ma$.

Conversely, if $\alpha$ is a $K$-linear transformation, $a\in R$, $m\in M$, then
$\alpha(\frac{a}{1}\circ m)=\frac{a}{1}\circ \alpha m$, i.e. $\alpha(am)=a\alpha(m)$. We have proved that every $K$-linear transformation is an right $R$-module homomorphism. \end{proof}

\begin{remark}\label{R:2}
The center $Z(R)$ of a weakly simple ring  $R$ is a domain.
\end{remark}
\begin{remark}\label{R:3}
For every right $R$-module $M$ the underlying group $M(+)$ is a discrete left topological   $({\rm End}\,(M_R),\mathcal{T}_{fin})$-module.
\end{remark}
Indeed, $T(m)(m)=0$ for every $m\in M$. Moreover,
\[
{\rm End}\,(M_R)\{0\}=\{0\},
\]
 so $M$ is a discrete left topological   $({\rm End}\,(M_R),\mathcal{T}_{fin})$-module.

\begin{theorem}\label{T:8}
Let $M_R$ be a module over a commutative ring $R$.

If the topological ring $({\rm End}\,(M_R),\mathcal{T}_{fin})$ is weakly simple, then:
\begin{itemize}
\item[{(i)}] $P=\{r\in R\mid Mr=0\}$ is a prime ideal of $R$.

\item[{(ii)}] $M$ is a vector space over the field $K$ of fractions of  $R/P$ and  the $R$-endomorphisms of $M$ are exactly the  $K$-linear transformations.

\end{itemize}

\noindent
Conversely,  if $M_R$ is an $R$-module and are satisfied (i) and (ii), then  $({\rm End}\,(M_R),\mathcal{T}_{fin})$ is a weakly simple topological ring.
\end{theorem}

\begin{proof}
$\Rightarrow:$
If  $({\rm End}\,(M_R),\mathcal{T}_{fin})$ is weakly simple, then  the mapping:
\begin{equation}\label{E:2}
\alpha_r: M \to M, \qquad  m\mapsto mr\qquad\quad(r\in R)
\end{equation}
is an $R$-module homomorphism and $\alpha_r\in Z$(= the center of ${\rm End}\,(M_R)$).

First we show that the part (i) holds.  Indeed, if $a,b\in R$ and $ab=0$, then $\alpha_a\alpha_b=0$ (see (\ref{E:2})). Thus $({\rm End}\,(M_R)\alpha_a)\cdot ({\rm End}\,(M_R)\alpha_b)=0$, so
\[
(\overline{{\rm End}\,(M_R)\alpha_a})\cdot (\overline{{\rm End}\,(M_R)\alpha_b})=0.
\]
Since ${\rm End}\,(M_R)$ is weakly simple, one of them,  say  $\overline{{\rm End}\,(M_R)\alpha_a}$, is zero.  This implies that $\alpha_a=0$, hence $a\in P$.

(ii) The structure of $R/P$-module  on $M$ is defined as follows:  if $r\in R$ and $m\in M$, then put $M(r+P)=mr$.

Note  that $M$ is a torsion-free right $R/P$-module. Assume  that  $m(r+P)=0$, where   $0\neq r+P\in R/P$ and  $0\neq m\in M$. Then $mr=0=\alpha_r(m)$ (see (\ref{E:2})).  Thus ${\rm End}\,(M_R)\alpha_r(m)=0$. It follows that  $\big(\overline{{\rm End}\,(M_R)\alpha_r}\big)(m)=0$ by  Remark \ref{R:3}. Since ${\rm End}\,(M_R)$ is weakly simple
\[
\overline{{\rm End}\,(M_R)\alpha_r}={\rm End}\,(M_R).
\]
We obtained that ${\rm End}\,(M_R)(m)=0$, so $m=0$, a contradiction.

 Under this convention  $R$-submodules are exactly $R/P$-submodules and $R$-endomorphisms are exactly
$R/P$-endomorphisms.

The module  $M$ is a divisible $R/P$-module. Indeed, if  $0\neq r+P\in R/P$, then  $0\not=M(r+P)=Mr$. Suppose that $Mr\neq M$. Consider
\[
I=\{\alpha\in E\vert \alpha(M)\subseteq Mr\}.
\]
Since $Mr$ is a fully invariant submodule, $I$ is a two-sided ideal of $({\rm End}\,(M_R),\mathcal{T}_{fin})$.

The ideal $I$ is closed. Indeed, let $\alpha\in \overline{I}$. If $m\in M$, then there exists $\beta\in I$ such that $\alpha-\beta\in T(m)$. Clearly,   $\alpha(m)=\beta(m)\in Mr$ and so $\alpha\in I$. We have proved that $I$ is closed.

Since $1_M\notin I$, $I=0$. It follows that    $\alpha_r=0$ (see (\ref{E:2})),  a contradiction.

The module $M$ has a structure of a right $K$-vector space and ${\rm End}\,(M_R)$ is exactly the ring of endomorphisms of $M$ by Lemma \ref{L:3}.

The converse follows from  Theorem \ref{T:6}.
\end{proof}

A characterization of completely simple topological ring ${\rm End}\,(M_R)$ is given by the following.

\begin{theorem}\label{T:9} Let $M_R$ be a module over a commutative ring $R.$
The topological  ring $({\rm End}\,(M_R), \mathcal{T}_{fin})$ is completely simple if and only  are satisfied the conditions (i) and (ii) of Theorem 8 and either
\begin{itemize}
\item[\rm(i)] $M$ is finite or
\item[\rm(ii)] $M$ is infinite and the dimension of $M$ over the field $K$  is infinite.
\end{itemize}
\end{theorem}

\begin{proof}
$\Rightarrow:$ According to Theorem \ref{T:8}, the ideal $P$ is prime   and the topology of  ${\rm End}\,(M_R)$ coincide with the finite topology of  ${\rm End}\,(M_K)$, where $K$ is the field of fractions of $R/P$. If $M$ is finite, we have the part (i).
Assume that $M$ is infinite. If $R/P$ is finite, then the dimension of $M$ over $K$ is infinite. Suppose that $R/P$ is infinite and $\dim _K(M)=n<\aleph_0$. Then $M$ is isomorphic to $M(n,K)$. Since $K$ is an infinite field, it admits a nondiscrete ring topology (see \cite{Hochster_Kiltinen}) and we obtain a contradiction because ${\rm End}\,(M_R)$ is a discrete ring. Consequently  $\dim _K(M)$ is infinite.

\noindent
$\Leftarrow$ This follows from Theorems \ref{T:6} and \ref{T:8}.\end{proof}

\begin{corollary}\label{C:4}
The topological ring $({\rm End}\,(A), \mathcal{T}_{fin})$ of an abelian group $A$ is completely  simple if and only  one of the following conditions holds:
\begin{itemize}
\item[\rm(i)] $A$ is a elementary abelian  $p$-group.
\item[\rm(ii)] $A$ is a divisible torsion-free group of infinite rank.
\end{itemize}
\end{corollary}

\newpage

\bibliographystyle{abbrv}
\bibliography{Bovdi_Ursul}
\end{document}